### v1

@@ 3.2 Projected Full-Attention (PFA)
Figure 3: The detailed workflow of PFA. The input tensor is first sent to Linear Projection of Spike Tensor (LPST) unit to generate three projections and split the three projections into corresponding vectors. In Attention Map Composing (AMC) unit, these vectors are composed into the final attention map through the inverse process of CP decomposition. The attention is fused with the input tensor to obtain the refined tensor by Hadamard product.
In this section, we first introduce the overall idea of PFA. Then we present the details of two crucial components of PFA, i.e., linear projection of spike tensor (LPST) and attention map composing (AMC).
The core of the attention mechanism is to give essential parts of the input data higher weights while ignoring less-related parts to some degree. This is particularly important for SNNs, especially for data collected by DVS cameras. DVS camera data often has more noise, which greatly affects network performance. Thanks to the feature of CP-decomposition, PFA can effectively suppress noise. In our proposed PFA, we first design the LPST module to project the higher-dimensional tensor (generally 3D) to three lower-dimensional tensors, i.e., 2D matrices. By using learnable parameters, LPST can selectively reserve essential information while ignoring irrelevant one. In AMC, we split the matrices into $r$ (called connecting factor here) vectors. Three vectors from three matrices of temporal, channel and spatial dimensions respectively are used to conduct the outer product to compose a part of the final attention map. The final attention map is the sum of these $r$ parts. We fuse this attention map with the origin input to obtain the refined tensor. For an input tensor with the shape of $H\times W\times C\times T$, the overall parameter amount is $C\times r+T\times r+k^{2}\times T\times r$, where $k$ is the convolution kernel size. Note that $r$ is limited to a small fixed value according to our theoretical analysis in Section 3.3, the parameter size grows linearly with $T$ and $C$. Figure 2 shows the parameter growth curve of PFA and compares several other techniques. The workflow of PFA is plotted in Figure 3. In what follows, we will give the details of LPST and AMC.
@@ 3.2.1 Linear Projection of Spike Tensor (LPST)
Tensor decomposition is a vital approach to represent a tensor and its latent features. Therefore it is natural for us to utilize tensor decomposition related techniques to explore the data characteristics. A typical technique is the CANDECOMP/PARAFAC (CP) decomposition. Whatever way we adopt, it is essential to generate a “projection” to extract significant components of the input tensor to yield an attention map later.
In this part, we need to construct three linear projections of the input data stream along spatial, channel, and temporal dimensions. In LPST, we apply two fully connected (FC) layers and a convolution layer to project the input tensor for simplicity. Note that here we simply use single-layer FCs and a convolution layer. What we do here is just a simple projection with learnable parameters rather than complex function fitting. In this way, we not only retain the required information, but also reduce the number of parameters.
For temporal dimension, we first squeeze the input tensor ($H\times W\times C\times T$) in the following way:
$Y^{T}_{(c,t)}=\frac{1}{HW}\sum\limits_{i=1}^{H}\sum\limits_{j=1}^{W}X_{(i,j,c,t)},$ (5)
where $Y^{T}_{(c,t)}$ denotes the output matrix, $H$ and $W$ denote the spatial sizes, $X_{(i,j,c,t)}$ denotes the input tensor. After the squeeze operation, the tensor is converted to a matrix in the shape of $C\times T$. Then, we apply an FC layer to reserve $r$ necessary information in each column. A sigmoid activation function is applied after the FC layer. Finally, we obtain a projection matrix in the shape of $r\times T$.
For channel dimension, the operation is similar to that of temporal dimension. We squeeze the input tensor by
$Y^{C}_{(c,t)}=\frac{1}{HW}\sum\limits_{i=1}^{H}\sum\limits_{j=1}^{W}X_{(i,j,c,t)},$ (6)
@@
For spatial dimension, the channel is squeezed by,
$Y^{S}_{(x,y,t)}=\frac{1}{C}\sum\limits_{k=1}^{C}X_{(x,y,k,t)}.$ (7)
Afterward, we employ a convolution to the squeezed tensor, treating $T$ as the input channel and $r$ as the output channel, and then reshape the output tensor from the convolution into a matrix in the shape of $HW\times r$. The sigmoid activation function is applied.
@@ 3.2.2 Attention Map Composing (AMC)
Once we obtain the three projections corresponding to three dimensions, a necessary step is to compose an attention map using the three projections. Here, we adopt the reverse process of the tensor CP decomposition. In the CP decomposition, a tensor $\mathcal{X}$ can be represented as:
$\mathcal{X}\approx\sum_{i=1}^{r}v_{si}\circ v_{ci}\circ v_{ti},$ (8)
where $v_{si}$, $v_{ci}$ and $v_{ti}$ denote the $i$-th vectors in the spatial, channel and temporal projection tensors respectively, and $\circ$ is the outer product of vectors. Especially, $r$ is the connecting factor, which can be viewed as the rank concept in the tensor CP decomposition (this process is plotted in Figure 3). It is worth noting that the tensor CP decomposition is just an approximate representation of tensors, which perfectly meets our demands. The accuracy of our representation of the input tensor can be controlled by r. That is the larger $r$ is, the more accurate the representation is, and vice versa, which perfectly realizes the attention mechanism we need. In the experiment part, we discuss the choice of connecting factor $r$ and its effect in details.
3.3 Quantitative Theoretical Analysis
@@
In this section, we first introduce an approach to estimate the rank of the input spike tensor. Following such analysis, we propose a principle to select the proper $r$ value, which is supported by our experiments.
To better understand the feature of PFA and facilitate later application, we follow the work by Acar et al. [1] to investigate the rank of the input data. Given that computing the CP decomposition of a tensor is recognized as an NP-hard problem [13; 22], direct calculation from the tensor becomes infeasible. Therefore, to address this challenge, we adopt a gradient descent technique to approximate the original tensor. We manually set the rank to multiple values, and consider the value with the best approximation as the rank of the tensor. To be more specific, an element of the tensor can be presented as:
@@
We treat it as an optimization problem, in which $a_{i,r},b_{j,r},c_{k,r}$ are considered as parameters. Then we optimize the $Loss$ in the following expression using the gradient descent approach:
$Loss=\frac{1}{2}\sum_{(i,j,k)\in\Omega}(x_{i,j,k}-\sum\limits_{r=1}^{R}a_{i,r}b_{j,r}c_{k,r})^{2}.$ (10)
Find the first order partial derivative of $Loss$ about the parameter ($a_{i,r}$ as an example here), namely:
$\frac{\partial Loss}{\partial a_{i,r}}=-\frac{1}{2}\cdot 2\cdot(x_{i,j,k}-\sum\limits_{r=1}^{R}a_{i,r}b_{j,r}c_{k,r})\cdot b_{j,r}c_{k,r}.$ (11)
Lastly, apply the gradient descent method to execute optimization in the following way:
$a_{i,r}^{\prime}\leftarrow a_{i,r}-\mu\cdot\frac{\partial Loss}{\partial a_{i,r}},$ (12)
where $\mu$ denotes the step size of the gradient descent. We take a piece of data from the dataset CIFAR10DVS as a toy example and discover its rank. In this case, we adopt $\mu=0.0001$ and iterate the process 1000 times to get the final optimized parameters. We utilize $\ell_{2}$ norm as the metric to judge the approximation effect. As shown in Figure 4, the error reduces as the rank increases, and starts to converge when the rank exceeds 30. This toy example shows that the rank of the input tensor is around 30, indicating that we need a rank exceeding 30 to precisely represent it using the CP decomposition technique. It should be noted that for the attention mechanism, representing the original tensor too accurately is not conducive to improving network performance. The core idea of the attention mechanism is to focus on the important parts while ignoring the secondaries. Representing raw tensors too precisely is counterproductive. Therefore, a good choice for $r$ should not be too large. This is a very useful conclusion. On the one hand, it tells us the range of finding $r$. On the other hand, it also ensures that PFA can obtain good results without excessive calculation.
@@ 4.1.1 Datasets
We have conducted experiments on both static datasets and dynamic datasets for object classification. We train and test PFA on a workstation equipped with one RTX 3090. For different datasets, the hyper-parameters are listed in Table 1. The summaries of datasets and augmentation involved in the experiment are listed below.
CIFAR 10/100 consist of 50k training images and 10k testing images with the size of 32 × 32 [24]. We use ResNet-19 for both CIFAR10 and CIFAR100. Random horizontal flips and crops are applied to the training images for augmentation. Moreover, Cutout is also used for augmentation which is the same as the Ref [8]
@@
NCALTECH-101 is also converted from the original version of Caltech-101 [32] with a slight change in object classes to avoid confusion. The NCaltech-101 consists of 100 object classes plus one background class. We apply the 9: 1 train-test split as CIFAR10DVS. And the VGGSNN is applied to neuromorphic datasets NCALTECH-101. Moreover, the augmentation is identical to CIFAR10DVS.
@@ 4.1.2 Loss function
We used the cross-entropy loss for the neuromorphic datasets CIFAR10DVS and NCALTECH-101. And in order to get a better performance in static datasets (CIFAR 10 / 100), we used a Temporal Efficient Training [8] loss function. It can be concluded as follows:
@@
where $T$ is the total time steps, $\mathcal{L}_{\mathrm{CE}}$ and $\mathcal{L}_{\mathrm{MSE}}$ denote the cross-entropy loss and mean square error. And $\boldsymbol{y}$ represents the target label. Moreover, $\mathbf{O}(t)$ and $\phi$ are the output and a constant used to regularize the membrane potential distribution. Additionally, we set $\phi=V_{th}$ in our experiments. In practice, we use a hyperparameter $\lambda$ to adjust the proportion of the regular term.
Figure 5: Schematic diagram of the network structure. On the left is the convolutional layer of the VGG network, and on the right is the convolutional layer of Resnet.
@@ 4.1.3 Network Architectures
To evaluate PFA, we integrate it into VGG-type [37] and ResNet [14] architectures, which are previously utilized in [12; 48; 8; 43]. Figure 5 presents the basic blocks involved in VGG and ResNet architectures. Five PFAs are inserted into a VGG-type network after each average pooling operation. In ResNet architecture, PFA is inserted before the residual connection in the last two layers. When splitting the event flow data into frames, we simply integrate the input in the time domain. Particularly, our work is based on the software framework SpikingJelly [10]. Taking notice that frames in CIFAR10-DVS and NCaltech-101 are entirely occupied by their content, we utilized a pooling layer as the first layer of the network to reduce memory consumption.
@@ 4.2 Comparison with Existing SOTA Works
To verify the effectiveness of our PFA, we compare it with several current best SNNs, including hybrid training and spike-based backpropagation. The results for the static image data and the dynamic data classification task are listed in Table 2. The results for the static image dataset are at time steps 2, 4, and 6.
For CIFAR10 and CIFAR100, PFA achieves higher accuracy than previous works. In particular, our PFA achieves superior performance on CIFAR10 and CIFAR100, with 1.2% and 4.4% improvements over the TET [8] method, respectively, using the same network structure and time steps. Moreover, our PFA surpasses the DSR [30] method while requiring 10$\times$ fewer time steps. Additionally, we demonstrate that our attention-based SNN outperforms the Spiking-ViT [47] backbone.
On the dynamic dataset CIFAR10DVS, our method outperforms the previous works with binary spikes by 3.1%, while using fewer time-steps. For the NCaltech101 dataset, we achieved 80.5% top-1 accuracy, achieving a performance far beyond previous work [48]. To further illustrate the advantages of our PFA, we show the comparison with previous works at different time steps. As shown in Figure 1, we compared PFA with Spike-based BP SNNs. Specifically, our PFA shows a significant advantage at shorter simulation lengths due to its more vital representation ability.
Dataset Method Type Architecture Time steps Accuracy CIFAR10/CIFAR100 SIDP [35]ICLR-2019 Hybrid training ResNet-20/VGGSNN 250/125 95.4/76.4 Diet [34]TNNLS-2021 Spike-based BP ResNet-20 10/5 92.5/64.0 tdBN [46]AAAI-2021 Spike-based BP ResNet-19 6 93.1/71.1 Dspike [27]NeurIPS-2021 Spike-based BP ResNet-18 6 94.2/74.2 NAS-SNN [21]ECCV-2022 Spike-based BP NAS 5 92.7/73.0 TET [8]ICLR-2022 Spike-based BP ResNet-19 6 94.5/74.7 DSR [30]CVPR-2022 Spike-based BP ResNet-19 20 95.4/78.5 Spikformer [47]ICLR-2023 Spike-based BP Spiking-ViT 4 95.5/78.2 PFA Spike-based BP ResNet-19 6 95.7/79.1 4 95.7/78.1 2 95.6/76.7 CIFAR10DVS LIAF-Net [40]TNNLS-2021 Spike-based BP LIAF-Net 10 70.4 tdBN [46]AAAI-2021 Spike-based BP ResNet-19 10 67.8 TA-SNN [42]ICCV-2021 Spike-based BP LIAF-Net 10 72.0 Dspike [27]NeurIPS-2021 Spike-based BP VGGSNN 10 75.4 TET [8]ICLR-2022 Spike-based BP VGGSNN 10 83.2 DSR [30]CVPR-2022 Spike-based BP VGGSNN 10 77.3 TCJA [48]arXiv-2022 Spike-based BP VGGSNN 10 80.7 Spikformer [47]ICLR-2023 Spike-based BP Spiking-ViT 16 80.9 PFA Spike-based BP VGGSNN 14 84.0 NCaltech-101 SALT [20]Neural Netw-2021 Spike-based BP VGGSNN 20 55.0 TCJA [48]arXiv-2022 Spike-based BP VGGSNN 14 78.5 PFA Spike-based BP VGGSNN 14 80.5

### v2

@@ 3.2 Projected Full-Attention (PFA)
Figure 3: The detailed workflow of PFA. The input tensor is first sent to Linear Projection of Spike Tensor (LPST) module to generate three projections and split the three projections into corresponding vectors. In Attention Map Composing (AMC) module, these vectors are composed into the final attention map through the inverse process of CP decomposition. The attention is fused with the input tensor to obtain the refined tensor by Hadamard product.
In this section, we will present an overview of our proposed PFA module and detailed descriptions of two sub-modules of PFA. PFA is composed of two sub-modules, Linear Projection of Spike Tensor (LPST) and Attention Map Composing (AMC). LPST, as its name suggests, produces projections from input tensor $\mathcal{X}$ for AMC in the form of matrices. The AMC module splits the matrices from LPST to vectors to compose the final attention map.
Figure 4: Comparison of accuracy between the vanilla VGG network and VGG network with PFA modules on training and validation sets. While the training set accuracy remains consistently high for both models, the validation set accuracy shows a significant improvement with the addition of PFA modules, suggesting its efficacy in mitigating overfitting.
The core of the attention mechanism is to give essential parts of the input data higher weights while ignoring less-related parts to some degree. This is particularly important for SNNs, especially for data collected by DVS cameras. Data from DVS camera often contains more noise, which greatly affects network performance. Thanks to the feature of CP-decomposition, PFA can effectively suppress noise. Also, as shown in Figure 4, the inexact representation of the input tensor can help reduce overfitting, which is very severe in SNNs. In our proposed PFA, we first design the LPST module to project the higher-order tensor (generally 3D) to three lower-order tensors, i.e., 2D matrices. Using learnable parameters, LPST can selectively reserve essential information while ignoring irrelevant ones. In AMC, we split the matrices into $R$ (called connecting factor here) vectors. Three vectors from three matrices of temporal, channel, and spatial dimensions respectively are used to conduct the outer product to compose a part of the final attention map. The final attention map is the sum of these $R$ parts. We fuse this attention map with the origin input to obtain the refined tensor.
The workflow of PFA is plotted in Figure 3. In what follows, we will give the details of LPST and AMC.
@@ 3.2.1 Linear Projection of Spike Tensor (LPST)
Tensor decomposition is a vital approach to represent a tensor and its latent features. Therefore, it is natural for us to utilize tensor-decomposition-related techniques to explore the data characteristics. A typical technique is the CP decomposition. Whatever way we adopt, it is essential to generate a ‘projection’ to extract significant components of the input tensor to yield an attention map later.
In this part, we need to construct three linear projections of the input data stream along spatial, channel, and temporal dimensions. In LPST, we apply two fully connected (FC) layers and a convolution layer to project the input tensor for simplicity. Note that here we simply use single-layer FCs and a convolution layer. What we do here is just a simple projection with learnable parameters rather than complex function fitting. In this way, we retain the required information and reduce the number of parameters.
For the temporal dimension, we first squeeze the input tensor ($H\times W\times C\times T$) in the following way:
$Y^{T}_{(c,t)}=\frac{1}{HW}\sum\limits_{i=1}^{H}\sum\limits_{j=1}^{W}\mathcal{X}_{(c,t,i,j)},$ (6)
where $Y^{T}_{(c,t)}$ denotes the output matrix, $H$ and $W$ denote the spatial sizes, $\mathcal{X}_{(i,j,c,t)}$ denotes the input tensor. After the squeeze operation, the tensor is converted to a matrix in the shape of $C\times T$. Then, we apply an FC layer to reserve $R$ necessary information in each column. A sigmoid activation function is applied after the FC layer. Finally, we obtain a projection matrix in the shape of $r\times T$.
For the channel dimension, the operation is similar to that of the temporal dimension. We squeeze the input tensor by:
$Y^{C}_{(c,t)}=\frac{1}{HW}\sum\limits_{i=1}^{H}\sum\limits_{j=1}^{W}\mathcal{X}_{(c,t,i,j)},$ (7)
@@
For spatial dimension, the channel is squeezed by:
$Y^{S}_{(t,x,y)}=\frac{1}{C}\sum\limits_{k=1}^{C}\mathcal{X}_{(k,t,x,y)}.$ (8)
Afterward, we employ a convolution to the squeezed tensor, treating $T$ as the input channel and $R$ as the output channel. Then, we reshape the output tensor from the convolution into a matrix in the shape of $HW\times r$. The sigmoid activation function is applied.
@@ 3.2.2 Attention Map Composing (AMC)
Once we obtain the three projections corresponding to three dimensions, a necessary step is to compose an attention map using the three projections. Here, we adopt the reverse process of the tensor CP decomposition. In the CP decomposition, a tensor $\mathcal{X}\in\mathcal{R}^{HW\times C\times T}$ can be represented as:
$\mathcal{A}_{s,c,t}=\sum_{r=1}^{R}U^{(HW)}_{s,r}\circ U^{(C)}_{c,r}\circ U^{(T)}_{t,r},$ (9)
where $U_{si}$, $U_{ci}$ and $U_{ti}$ denote the $i$-th vectors in the spatial, channel, and temporal projection tensors respectively, $\mathcal{A}$ is the attention map, and $\circ$ is the outer product of vectors. In particular, $R$, the connecting factor, can be viewed as the rank concept in the tensor CP decomposition (this process is plotted in Figure 3). It is worth noting that the tensor CP decomposition is just an approximate representation of tensors, which perfectly meets our demands. $R$ can control the accuracy of our representation of the input tensor. That is, the larger $R$ is, the more accurate the representation is, and vice versa, which perfectly realizes the attention mechanism we need. In the experiment part, we discuss the choice of connecting factor $R$ and its effect in detail.
3.3 Parameter and Computational Cost Analysis
This section presents an analysis of the parameter and computational cost of PFA. PFA has an advantage in terms of parameter quantity. For an input tensor with the shape of $H\times W\times C\times T$, the overall parameter amount is:
$\underbrace{C\times R}_{\text{MLP of }Y^{T}}+\underbrace{T\times R}_{\text{MLPof }Y^{C}}+\underbrace{k^{2}\times T\times R}_{\text{Conv of }Y^{S}},$ (10)
where $k$ is the convolution kernel size. Note that $R$ is limited to a small fixed value according to our theoretical analysis in Section 3.4, and the parameter size grows linearly with $T$ and $C$. Figure 2 shows the parameter growth curve of PFA and compares several other techniques.
PFA is also lightweight in computational burden. The time complexity of PFA can be calculated in the following way:
$\underbrace{3HWTC}_{\text{Obtain }Y^{T},Y^{C},Y^{S}}+\underbrace{2TCR}_{\text{MLP of }Y^{T},Y^{C}}+\underbrace{HWk^{2}TR}_{\text{Conv of }Y^{S}}+\underbrace{RHWTC}_{\text{AMC}}.$ (11)
It is worth noting that a standard convolution operation typically entails a computational burden of $HWk^{2}TC_{in}C_{out}$. In this context, given that $R$ is not large, a PFA module incurs a computational load that is smaller than that of a convolution operation since $C$ is usually in the range of several hundred.
3.4 Theoretical Analysis on $R$
@@
In light of the fact that CP decomposition provides only an approximate representation of a tensor, and the quality of this approximation relies on the chosen factor, denoted as $R$, it is essential to establish a clear guiding principle for determining the optimal value of $R$. This section delves into our approach for gaining insights into the characteristics of PFA and streamlining its future applications.
Our exploration begins by adopting the methodology put forth by Acar et al. [1] to investigate the rank of the input data. However, due to the recognized NP-hard nature of computing the CP decomposition of a tensor [13; 24], a direct calculation from the tensor becomes impractical. To surmount this challenge, we employ a gradient descent technique to approximate the original tensor by systematically experimenting with different rank values, ultimately selecting the one that yields the best approximation for the tensor. To provide further clarity, consider an element of the tensor represented as:
@@
We treat this as an optimization problem, with $a_{i,r}$, $b_{j,r}$, and $c_{k,r}$ considered as parameters. We then optimize the loss function defined as:
$\mathcal{L}=\frac{1}{2}\sum_{(i,j,k)\in\Omega}(x_{i,j,k}-\sum\limits_{r=1}^{R}a_{i,r}b_{j,r}c_{k,r})^{2}.$ (13)
To perform optimization, we compute the first-order partial derivative of the loss with respect to a parameter (e.g., $a_{i,r}$), given by:
$\frac{\partial\mathcal{L}}{\partial a_{i,r}}=-\frac{1}{2}\cdot 2\cdot(x_{i,j,k}-\sum\limits_{r=1}^{R}a_{i,r}b_{j,r}c_{k,r})\cdot b_{j,r}c_{k,r}.$ (14)
Finally, we apply the gradient descent method for optimization:
$a_{i,r}^{\prime}\leftarrow a_{i,r}-\mu\cdot\frac{\partial\mathcal{L}}{\partial a_{i,r}},$ (15)
Here, $\mu$ represents the step size used in the gradient descent process. This approach allows us to effectively estimate the rank of the tensor while circumventing the computational complexities associated with CP decomposition.
We take a piece of data from the dataset CIFAR10DVS as a toy example and discover its rank. In this case, we adopt $\mu=0.0001$ and iterate the process 1000 times to get the final optimized parameters. We utilize $\ell_{2}$ norm as the metric to judge the approximation effect. As shown in Figure 5, the error reduces as the rank increases, and starts to converge when the rank exceeds 30. This toy example shows that the rank of the input tensor is around 30, indicating that we need a rank exceeding 30 to represent it using CP decomposition precisely. It should be noted that accurately representing the original tensor for the attention mechanism is not conducive to improving network performance. The core idea of the attention mechanism is to focus on the important parts while ignoring the secondaries. Representing raw tensors too precisely is counterproductive. On the one hand, overly precise characterization of input tensors is detrimental to neglecting unimportant parts. On the other hand, it may lead to severe overfitting. As shown in Figure 4, the vanilla VGG network exhibits severe overfitting, achieving nearly 100% accuracy on the training set but less than 80% on the validation set. However, upon incorporating the PFA modules, there is not a substantial improvement in training set accuracy. Nevertheless, there is a significant increase in validation set accuracy. This observation underscores the benefit of utilizing less precise tensor representations with PFA to mitigate overfitting. Therefore, a good choice for $R$ should not be too large. This is a very useful conclusion. On the one hand, it tells us the range of finding $R$. On the other hand, it also ensures that PFA can obtain good results without excessive calculation.
@@ 4.1.1 Datasets
We have conducted experiments on both static and dynamic datasets for object classification. We train and validate PFA on a workstation equipped with one RTX 3090. For different datasets, the hyper-parameters are listed in Table 1. The summaries of datasets and augmentation involved in the experiment are listed below.
CIFAR 10/100 consist of 50k training images and 10k testing images with the size of $32\times 32$ [30]. We use ResNet-19 for both CIFAR10 and CIFAR100. Random horizontal flips and crops are applied to the training images for augmentation. Moreover, cutout is also used for augmentation which is the same as [8]
@@
NCALTECH-101 is also converted from the original version of Caltech-101 [40] with a slight change in object classes to avoid confusion. The NCaltech-101 consists of 100 object classes plus one background class. We apply the 9: 1 train-validation split as CIFAR10DVS. The VGGSNN is applied to neuromorphic datasets NCALTECH-101. Moreover, the augmentation is identical to CIFAR10DVS. Fashion-MNIST is a more demanding successor of the famous MNIST dataset. Fashion-MNIST contains 70,000 static grayscale images of 10 different categories of clothing, each image being of size 28x28 pixels. We utilize the simple example network from the Spikingjelly framework, which comprises two convolution layers and a fully-connected layer, to evaluate PFA on this dataset.
@@ 4.1.2 Loss function
We used the cross-entropy loss for the neuromorphic datasets CIFAR10DVS and NCALTECH-101. In order to get a better performance in static datasets (CIFAR 10 / 100), we used a Temporal Efficient Training [8] loss function. It can be concluded as follows:
@@
where $T$ is the total time steps, $\mathcal{L}_{\mathrm{CE}}$ and $\mathcal{L}_{\mathrm{MSE}}$ denote the cross-entropy loss and mean square error. And $\boldsymbol{y}$ represents the target label. Moreover, $\mathbf{O}(t)$ and $\phi$ are the output and a constant used to regularize the membrane potential distribution. Additionally, we set $\phi=V_{th}$ in our experiments. In practice, we use a hyper-parameter $\lambda$ to adjust the proportion of the regular term.
Figure 6: Schematic diagram of the network structure. On the left is the convolution layer of the VGG network, and on the right is the convolution layer of Resnet.
@@ 4.1.3 Network Architectures
To evaluate PFA, we integrate it into VGG-type [47] and ResNet [14] architectures, which are previously utilized in [12; 70; 8; 63]. Figure 6 presents the basic blocks involved in VGG and ResNet architectures. Five PFA modules are inserted into a VGG-type network after each average pooling operation. In the ResNet architecture, PFA is inserted before the residual connection in the last two layers. When splitting the event flow data into frames, we simply integrate the input in the time domain. Particularly, our work is based on the software framework SpikingJelly [10]. Taking notice that frames in CIFAR10-DVS and NCaltech-101 are entirely occupied by their content, we utilized a pooling layer as the first layer of the network to reduce memory consumption.
@@ 4.2 Comparison with Existing SOTA Works
Dataset Method Type Architecture Time steps Accuracy CIFAR10/CIFAR100 SIDP  [44]ICLR-2019 Hybrid training ResNet-20/VGGSNN 250/125 95.4/76.4 Diet [43]TNNLS-2021 Spike-based BP ResNet-20 10/5 92.5/64.0 tdBN [68]AAAI-2021 Spike-based BP ResNet-19 6 93.1/71.1 Dspike [34]NeurIPS-2021 Spike-based BP ResNet-18 6 94.2/74.2 NAS-SNN [23]ECCV-2022 Spike-based BP NAS 5 92.7/73.0 TET [8]ICLR-2022 Spike-based BP ResNet-19 6 94.5/74.7 DSR [37]CVPR-2022 Spike-based BP ResNet-19 20 95.4/78.5 Spikformer [69]ICLR-2023 Spike-based BP Spiking-ViT 4 95.5/78.2 PFA Spike-based BP ResNet-19 6 95.7/79.1 4 95.7/78.1 2 95.6/76.7 CIFAR10DVS LIAF-Net  [57]TNNLS-2021 Spike-based BP LIAF-Net 10 70.4 tdBN [68]AAAI-2021 Spike-based BP ResNet-19 10 67.8 TA-SNN  [62]ICCV-2021 Spike-based BP LIAF-Net 10 72.0 Dspike [34]NeurIPS-2021 Spike-based BP VGGSNN 10 75.4 TET [8]ICLR-2022 Spike-based BP VGGSNN 10 83.2 DSR  [37]CVPR-2022 Spike-based BP VGGSNN 10 77.3 TCJA  [70]TNNLS-2024 Spike-based BP VGGSNN 10 80.7 Spikformer [69]ICLR-2023 Spike-based BP Spiking-ViT 16 80.9 PFA Spike-based BP VGGSNN 14 84.0 NCaltech-101 SALT  [22]Neural Netw-2021 Spike-based BP VGGSNN 20 55.0 TCJA  [70]TNNLS-2024 Spike-based BP VGGSNN 14 78.5 PFA Spike-based BP VGGSNN 14 80.5
@@
To validate the effectiveness of our PFA, we conducted a comparative analysis with several state-of-the-art SNNs, including those employing hybrid training and spike-based backpropagation. The results for both static image data and dynamic data classification tasks are summarized in Table 2. Specifically, the results for the static image dataset are presented at time steps 2, 4, and 6.
For CIFAR10 and CIFAR100 datasets, our PFA outperforms prior works in terms of accuracy. Notably, PFA achieves superior performance on CIFAR10 and CIFAR100, with respective improvements of 1.2% and 4.4% over the TET method [8] which utilizes the same network architecture and time steps. Furthermore, PFA surpasses the DSR method [37] while requiring only 10$\times$ fewer time steps. Additionally, our attention-based SNN demonstrates better performance than the Spiking-ViT backbone [69].
Method Time Steps Accuracy
ST-RSBP[66]${}^{\text{NIPS-2019}}$ 400 90.1
LISNN[6]${}^{\text{IJCAI-2020}}$ 20 92.1
PLIF[12]${}^{\text{ICCV-2021}}$ 8 94.4
TCJA[70]${}^{\text{TNNLS-2024}}$ 8 94.8
PFA (Ours) 8 94.5
Table 3: Accuracy on Fashion-MNIST[59] dataset.
For the Fashion-MNIST dataset, as depicted in Table 3, while the accuracy of our proposed PFA falls slightly short of TCJA [70], it still surpasses previous methodologies utilizing longer time steps [66; 6] or incorporating learnable spike thresholds [12]. We hypothesize that the relatively diminutive image size of the Fashion-MNIST dataset curtails the efficacy of spatial attention. This conjecture aligns with the findings from our ablation experiments, indicating that spatial attention plays a pivotal role in determining PFA performance.
For the dynamic dataset CIFAR10DVS, our method outperforms previous approaches using binary spikes by 3.1%, even with fewer time steps. For the NCaltech101 dataset, we achieved an impressive 80.5% top-1 accuracy, far surpassing the performance of prior work [70].
To further illustrate the advantages of our PFA, we compare it with previous works at various time steps. As depicted in Figure 1, we compare PFA with Spike-based BP SNNs. Notably, our PFA exhibits a significant advantage at shorter simulation time steps, owing to its superior representation capability.
We observe from the perspective of simulated time steps that PFA can achieve satisfactory performance even with a smaller number of simulated time steps. For the CIFAR-10 dataset, simply utilizing 2 simulated time steps with PFA can surpass the performance of previous works, including transformer-like structures such as Spikeformer. On the CIFAR-100 dataset, compared to the top-performing Spikeformer, PFA lags by only 0.1% accuracy with 4 time steps. On DVS datasets like CIFAR-10DVS and NCaltech, employing 14 simulated time steps, PFA significantly outperforms previous methods. On CIFAR10-DVS, PFA surpasses Spikeformer by 3.1%, and compared to the TET [8] method with a similar structure, it achieves a higher accuracy by 0.8%. On the NCaltech-101 dataset, PFA outperforms the previous best TCJA method by 2%.
In summary, PFA offers substantial advantages over previous methods. When applied to attention modules like TCJA [70], PFA significantly enhances resistance to overfitting, leading to improved accuracy on the validation set. Although Spikeformer, which is based on the Spike ViT architecture, performs better than many CNN-based techniques, an especially interesting breakthrough is the incorporation of PFA into traditional CNN architectures like VGG, which can even outperform Spikeformer’s [69] performance.
4.3 Experiments on Image Generation Tasks
Figure 7: Generated images of ANN VAE, FSVAE[21], and our PFA on CIFAR10 [29] dataset.
[…]
MNIST[7] ANN[21]AAAI-2022 5.95 112.5 17.09
FSVAE[21]AAAI-2022 6.21 97.06 35.54
TCJA[70]TNNLS-2024 6.45 100.8 19.39
PFA (Ours) 6.97 96.82 15.12
Fashion- MNIST[59] ANN[21]AAAI-2022 4.58 123.7 18.08
FSVAE[21]AAAI-2022 4.55 90.12 15.75
TCJA[70]TNNLS-2024 5.61 93.41 12.46
PFA (Ours) 5.35 97.3 11.97
CIFAR10[29] ANN[21]AAAI-2022 2.59 229.6 196.9
FSVAE[21]AAAI-2022 2.94 175.5 133.9
TCJA[70]TNNLS-2024 3.73 170.1 100.4
PFA (Ours) 3.84 166.4 92.83
Table 4: Comparison with existing methods on image generation tasks.
To further validate the efficacy of PFA, we conduct image generation tasks on MNIST [7], Fashion-MNIST [59], and CIFAR10 [29] using fully spike variation autoencoders [21] (FSVAE) . We integrated our proposed PFA as the first layer of FSVAE while maintaining other settings identical to the original network. The experimental results are outlined in Table 4. Notably, PFA exhibits superior performance metrics on both MNIST [7] and CIFAR10 [29] datasets. Additionally, we visually depict the generated images in Figure 7, further corroborating the enhancements facilitated by PFA.